# Bridging the Divide: Unraveling the Knowledge Gap in Data Visualization Research and Practice

### v2

@@ III-C Guideline Landscape
We endeavored to code data types using the well-established framework in the research community, which includes categorical, numerical, temporal, and geospatial data types. However, we observed that design guidelines rarely specify such data types (Figure 1), often resulting in unreliable codes. Such data types could often be inferred from the chart types, however. We used Amar et al.[66]’s taxonomy of low-level analytic tasks to code task types. A significant portion of the 150 guidelines (38%), where task types were available, centered around the basic tasks of compare (18%) and retrieve value (7%); compare is a combination of retrieve value and compute derived value. Other task types observed include correlate (5%), characterize distribution (3%), and sort (2%).
@@ III-D Guideline Templates
While inspecting guideline sources during the coding process, we observed recurring patterns in the guideline structures. We categorize them into three templates: problem-solving, enhancement, alternative suggestion. We synthesized these patterns to derive a holistic guideline template (Figure 2): What/if {Context}s, {Approach* + (Alternative)}, because of {Problem} + for {Purpose}.
@@
Alternative suggestion guidelines are similar to problem-solving guidelines, but they warrent separate consideration given their prevalence in practice. Most of the guidelines that have user alternative chart (24%) and add different encoding (8%) approaches belong to this category; for instance, when you need more than six categorical colors, try alternative visual encoding such as position.
@@
The existing guideline description lacked a cohesive structure; however, these templates provide a well-defined framework for effectively conveying actionable guidance to designers. We used these templates to generate concise one-sentence descriptions for each guideline, further streamlining our open coding process.
@@ III-E Guideline and Empirical Study Alignment
Among the 390 collected guidelines, 109 (28%) were mapped to empirical studies. Among the 235 empirical studies, 94 were linked to the guidelines. In other words, each guideline was often mapped to more than one empirical study ($avg$ / guideline = 2.00, $std$=1.62), while an empirical study was also mapped to more than one guideline ($avg$ / empirical study = 2.15, $std$=1.72). Many of the highly mapped guidelines and empirical studies were about well-known graphical perception issues such as regarding pie charts, embellishments, color maps, and scales & axes.
Among the total 219 guideline and empirical study mappings, the majority (49%) were found to be supported by corresponding empirical studies. However, we also observed mixed support (32%) or even conflicting results (21%) from empirical studies. For example, a guideline suggests avoiding the use of 3D charts because they can skew perception. However, a study found that 3D heatmaps were superior to 2D heatmaps for reading and comparing single data items, but 2D heatmaps were better for overview tasks [67]. Similarly, some guidelines recommend minimalist chart designs without embellishments to avoid distracting from the data, but studies suggest that visual embellishments such as icons can improve long-term recall and interpretation accuracy [68]. Similarly, other guidelines recommend avoiding rainbow colormaps, while several empirical studies contradict this guideline by demonstrating the value of multi-hue maps on continuous data for class discriminability [69].
@@
Mapped empirical study topics Unmapped empirical study topics
1 point, bar, depict, rank, simple 1 color, bias, background, work, interaction
2 different, popout, range, conduct, effect 2 uncertainty, link, flow, different, text
3 ratio, trend, measure, similarity, aspect 3 different, encode, mark, value, position
4 effect, caption, convey, form, color 4 accuracy, time, compare, different, overlay
5 find, correlation, model, encode, embellishments 5 title, encode, value, pair, compare
6 line, accuracy, people, alternative, qualitative 6 time, different, evaluate, plot, series
7 color, different, encode, perception, map 7 graph, cluster, present, base, model
8 forecast, structure, size, trust, scatterplots 8 present, interactive, effect, people, type
9 wrap, compare, model, scale, value 9 color, edge, base, perception, present
10 color, size, concepts, mark, time 10 size, network, base, large, node
TABLE I: Comparative analysis of mapped and unmapped empirical studies through topic modeling conducted on the abstracts of studies. While many topics are broadly similar across the two groups, specific themes like networks, uncertainty, interaction, and biases were unique to these unmapped studies.
While we did not codify empirical studies, we conducted a brief analysis on unmapped ones to identify knowledge possibly unaddressed in practice. We utilized Latent Dirichlet Allocation, setting the number of topics to 10, for analyzing paper abstracts (Table I). Generally, these topics covered similar perception and cognitive issues, but the unmapped studies revealed several topics such as network, graph layout, cluster, interaction, uncertainty, and bias, suggesting areas that might be underrepresented in existing practice-oriented guidelines.
@@ III-F Takeaways
The analysis of guidelines and their alignment with empirical studies has yielded valuable insights into the design knowledge currently available and its empirical backing. We found that most guidelines primarily address basic charts, highlighting a gap in advanced visualizations like parallel coordinates and network diagrams. Interestingly, empirical evidence sometimes contradicted these guidelines, indicating the presence of mixed and conflicting knowledge. While variation is expected due to different contexts, it is potentially problematic when practitioners are not made aware of these nuances. Likewise, many guidelines lacked specific information on data and task types, critical for understanding the applicability of guidelines. Finally, the presence of unmapped empirical studies suggests opportunities for refining or creating new guidelines. Similarly, guidelines that remain unmapped point to areas where further empirical research is needed to develop solid evidence.
@@ IV-A Survey Recruitment & Procedure
Our recruitment methodology varied slightly between practitioners and researchers. For practitioners, our recruitment targeted individuals engaged in data visualization, regardless of their specific job titles and responsibilities, similar to prior work in data visualization practice [14]. This recruitment was based on the premise that knowledge of data visualization design is essential for anyone who creates visualizations. We reached out to members of the Data Visualization Society’s Slack workspace [13], focusing on individuals who had shown recent engagement through activities such as posting messages in public channels. Our approach involved the use of direct messaging, intended to improve the overall participation rate. Additionally, we endeavored to expand our reach through postings on the Storytelling with Data forum [70] and direct messages to visualization practitioners on Twitter, although these efforts were not as successful as the initial approach.
@@
We distributed the surveys online, and participants took approximately 6-10 minutes to complete each. The surveys contained similar questions, categorized into sections on background, awareness, opinions on visualization design guidelines, and perceptions and expectations of empirical research in visualization (see example questions in Figure 4 and Figure 5). We presented participants with examples of misleading visuals from WTF Viz [71] to gauge their awareness of such charts. We also included a screenshot of a comparative experiment between bar and pie charts to assess familiarity with empirical research. The full survey questionnaires can be found in the supplementary material. To compensate participants, we offered a \$25 raffle reward for every 25 participants who completed the survey.
@@ IV-C Survey Results
Both stakeholders agreed that guidelines could be beneficial for creating better visualizations (Practitioners: 27%, Researchers: 28%), improving confidence in design decisions (P: 20%, R: 23%), and facilitating better communication with stakeholders (P: 20%, R: 22%)). However, researchers weighed the usefulness of guidelines on increasing visualization design knowledge less (15%) than practitioners (24%). In the Other comment section, several practitioners mentioned additional benefits of guidelines: serving as a “standard of practice”; helping avoid “misleading, ambiguous”, and “deceitful” visualization designs; increasing “user understanding” and “trust”, producing “consistent user experience”; “facilitating design decision making”. Researchers’ comments were similar, while they additionally mentioned that guidelines provide a “scaffold” or “starting point” for learners.
@@ IV-E Interview Participants
TABLE II: Demographic characteristics of practitioner and researcher interviewees
@@ IV-F Interview Analysis
All interviews were conducted within a 30 to 45-minute timeframe, and each was transcribed using an automatic transcription service. The transcripts were then manually checked for errors. During the initial coding process, a research assistant pulled related quotes from the transcripts and coded them in a spreadsheet using the interview questions as the initial set of codes. The lead researcher reviewed and refined the codes. They repeated this process for approximately one-third of the interviews until arriving at a consistent set of codes. The research assistant then applied the final codes to the remaining interviews, which were comprehensively reviewed by the lead researcher. Overall, we categorized the quotes into eleven themes: five focused on experience and perception, challenges related to guidelines, two themes concerning perception and accessibility of empirical studies, four themes addressing the perception of the knowledge gap, strategies to mitigate this gap, and expectations within the visualization communities.
@@ IV-G2 Examples of Visualization Guidelines
Researchers echoed practitioners in some areas, reiterating guidelines on zero-baselines (R1, R10) and pie charts (R7). But they also shared more specific design knowledge such as rainbow color maps (R2, R6, R9), length vs angle vs area (R1, R2, R10), comparison along a common-scale vs unaligned-scale (R1, R8), 3D charts (R4), and banking to 45 degrees (R9). Some of these guidelines were neither found on our guideline analysis nor discussed by practitioners. Researchers were often critical in their discussions, noting exceptions to general guidelines and pointing out vagueness in some common advice. For instance, R6 said, “there are certain cases when with the scientific data it makes sense to actually utilize [rainbow colormaps]”, while R7 said, “One vis design guideline that we hear a lot about is don’t use pie charts, which I would take to be a bit vague.”
@@ IV-I4 Ideas and Methods for Alleviating the Gap
V Discussion
@@
First of all, there appears to be a need to establish consistent vocabularies for describing data visualizations, including chart names and task types. This issue was observed during our examination of guidelines and was reiterated in interviews with practitioners; e.g., there is often a lack of clear distinction between task types, chart types, and data types in the categorization of visualizations [46, 72, 47]. While researchers may require a more granular vocabulary, such as the task taxonomy proposed by Brehmer and Munzner [28], it must be tailored to the level of practical needs and contexts.
A critical issue we observed is inconsistent design knowledge available in practice, as well as authoritative guidelines that do not acknowledge contextual factors. Interviewees often recognized this issue and suggested the need for improving current design guidelines with empirical evidence details. Empirical studies often contain mixed and conflicting findings, reflecting the diverse contexts in which these studies are conducted. However, the challenge lies in effectively communicating these complex nuances to practitioners.
@@
We also observed how practitioners appreciate intuition and experience in design practice. We need to consider how to incorporate such experiential knowledge. VisGuides [42] is an example of a platform where practitioners can share their examples for discussion and feedback. We can also imagine connecting existing examples to guidelines to create a triadic model of guidelines, empirical evidence, and relevant examples. The examples can demonstrate how people applied or deviated from the guidelines and present their success and failure cases. Furthermore, these examples might inspire researchers to develop future practice-oriented studies, thereby building a feedback and feedforward loop.
@@
Participants also shared the difficulty of finding relevant guidelines. Therefore, annotating the guidelines with detailed metadata, such as tasks, chart types, or other contextual factors, would be necessary to facilitate efficient navigation of the landscape of visualization design knowledge. Current efforts by practitioners [46, 73, 47] provide hints on how this idea can be realized. For example, a decision tree can be used to find a chart based on data types or task goals. However, as mentioned earlier, practitioners’ taxonomy is currently not consistent and streamlined compared to the research side. To alleviate this issue, we could utilize templates to create more structured and easily consumable guidelines, as demonstrated in Figure 2. These templates can be transformed into computational constraints to enhance a visualization recommendation system [74, 75].
@@
While building such a design knowledge base is feasible given the relatively young field of visualization the substantial ongoing challenge lies in maintaining its relevance and functionality over time. We believe that it will require collective efforts from both practitioners and researchers. Even if there have been efforts for both stakeholders to meet and collaborate, we have not yet seen a cohesive community that brings both together. For example, IEEE VIS primarily caters to researchers, while the Data Visualization Society focuses on practitioners. Given the challenges of creating a new community site [76], a viable alternative is to utilize existing communities to foster collaborations.
@@
Practitioners often write guidelines to promote their commercial tools (e.g., DataWrapper [54]) or attract clients (e.g., Data Viz Project [47]), similar to how researchers write blog articles for recognition. Yet, interviews highlighted a lack of incentives as a major issue. Addressing this, we should consider providing both intrinsic and extrinsic motivation to support the knowledge base’s maintenance and development. Ko’s market-based incentive model [77], offering tokens for contributions, presents an intriguing solution. Additionally, harnessing recent advancements in large language models could offer intelligent assistance in converting empirical results into practical guidelines [78].
@@
The field of data visualization is not alone in facing a knowledge gap between research and practice; this is a common challenge across many disciplines. In medicine, evidence-based practices integrate clinical expertise with the best available research evidence [11]. In policy-making, there is an emphasis on translating research findings into actionable policies and practices [9]. In education, it is common to use educational research to improve teaching and learning practices [10]. Learning from these established fields can also inform how the visualization field should progress to promote evidence-based design practice, such as developing knowledge brokering roles [15] and investing collaborative research-practice partnerships [79].
VI Conclusion

### v3

@@ III-C Guideline Landscape
We endeavored to code data types using the well-established framework in the research community, which includes categorical, numerical, temporal, and geospatial data types. However, we observed that design guidelines rarely specify such data types (Figure 1), often resulting in unreliable codes. Such data types could often be inferred from the chart types, however. We used Amar et al.[amar2005low]’s taxonomy of low-level analytic tasks to code task types. A significant portion of the 150 guidelines (38%), where task types were available, centered around the basic tasks of compare (18%) and retrieve value (7%); compare is a combination of retrieve value and compute derived value. Other task types observed include correlate (5%), characterize distribution (3%), and sort (2%).
@@ III-D Guideline Templates
While inspecting guideline sources during the coding process, we observed recurring patterns in the guideline structures. We categorized and developed them into three templates: problem-solving, enhancement, alternative suggestion. We synthesized these patterns to derive a holistic guideline template (Figure 2): What/if {Context}s, {Approach* + (Alternative)}, because of {Problem} + for {Purpose}.
@@
Alternative suggestion guidelines are similar to problem-solving guidelines, but they warrant separate consideration given their prevalence in practice. Most of the guidelines that have use alternative chart (24%) and add different encoding (8%) approaches belong to this category; for instance, when you need more than six categorical colors, try alternative visual encoding such as position.
@@
The existing guideline description lacked a cohesive structure. These templates can provide a well-defined framework for effectively conveying actionable guidance to designers, akin to efforts in medicine for developing models of guideline representation [ohno1998guideline]. We used these templates to generate concise one-sentence descriptions for each guideline, further streamlining our open coding process.
@@ III-E Guideline and Empirical Study Alignment
Among the 390 collected guidelines, 109 (28%) were mapped to empirical studies. Among the 235 empirical studies, 94 were linked to the guidelines. In other words, each guideline was often mapped to more than one empirical study (avgavg / guideline = 2.00, stdstd=1.62), while an empirical study was also mapped to more than one guideline (avgavg / empirical study = 2.15, stdstd=1.72). Many of the highly mapped guidelines and empirical studies were about well-known graphical perception issues such as regarding pie charts, embellishments, color maps, and scales & axes.
Among the total 219 guideline and empirical study mappings, the majority (49%) were found to be supported by corresponding empirical studies. However, we also observed mixed support (32%) or even conflicting results (21%) from empirical studies. For example, a guideline suggests avoiding the use of 3D charts because they can skew perception. However, a study found that 3D heatmaps were superior to 2D heatmaps for reading and comparing single data items, but 2D heatmaps were better for overview tasks [kraus2020assessing]. Similarly, some guidelines recommend minimalist chart designs without embellishments to avoid distracting from the data, but studies suggest that visual embellishments such as icons can improve long-term recall and interpretation accuracy [haroz2015isotype]. Similarly, other guidelines recommend avoiding rainbow colormaps, while several empirical studies contradict this guideline by demonstrating the value of multi-hue maps on continuous data for class discriminability [liu2018somewhere].
@@ III-F Takeaways
The analysis of guidelines and their alignment with empirical studies has yielded valuable insights into design guidelines currently available and its empirical backing. We found that most guidelines primarily address basic charts, highlighting a gap in advanced visualizations like parallel coordinates and network diagrams. Interestingly, empirical evidence sometimes contradicted these guidelines, indicating the presence of mixed and conflicting knowledge. While variation is expected due to different contexts, it is potentially problematic when practitioners are not made aware of these nuances. Likewise, many guidelines lacked specific information on data and task types, critical for understanding the applicability of guidelines. Finally, the presence of unmapped empirical studies suggests opportunities for refining or creating new guidelines. Similarly, guidelines that remain unmapped point to areas where further empirical research is needed to develop solid evidence.
III-G Limitations
While the core messages we intend to convey remain intact, the precise mapping statistics are transient, as they can easily shift with the emergence of new guidelines and the choices made regarding data sources and inclusion/exclusion criteria. As mentioned earlier, guidelines found in the wild are highly heterogeneous and difficult to categorize precisely due to the varied language, structure, and data quality used in different sources (e.g., the use of very generic language). We applied a reasonable rationale and perspective when selecting guidelines for our analysis (Section III-A1), aiming to highlight meaningful gaps to address. However, we discourage draw any significant conclusions from the absolute or relative percentage differences.
Moreover, our inclusion of less common guidelines may overstate the perceived knowledge gap, although this inclusion is important as it highlights underexplored areas of design knowledge. In addition, our alignment analysis was limited to empirical studies within the visualization literature, and it is possible that some unmatched guidelines could be supported by evidence from broader fields, such as cognitive psychology or art theory, which were outside the scope of this work.
@@ IV-A Survey Recruitment & Procedure
Our recruitment methodology varied slightly between practitioners and researchers. For practitioners, our recruitment targeted individuals engaged in data visualization, regardless of their specific job titles and responsibilities, similar to prior work in data visualization practice [parsons2021understanding]. This recruitment was based on the premise that knowledge of data visualization design is essential for anyone who creates visualizations. We reached out to members of the Data Visualization Society’s Slack workspace [datavissociety], focusing on individuals who had shown recent engagement through activities such as posting messages in public channels. Our approach involved the use of direct messaging, intended to improve the overall participation rate. Additionally, we endeavored to expand our reach through postings on the Storytelling with Data forum [storytellingwithdata] and direct messages to visualization practitioners on Twitter, although these efforts were not as successful as the initial approach.
@@
We distributed the surveys online, and participants took approximately 6-10 minutes to complete each. The surveys contained similar questions, categorized into sections on background, awareness, opinions on visualization design guidelines, and perceptions and expectations of empirical research in visualization (see example questions in Figure 4 and Figure 5). We presented participants with examples of misleading visuals from WTF Viz [wtfviz] to gauge their awareness of such charts. We also included a screenshot of a comparative experiment between bar and pie charts to assess familiarity with empirical research. The full survey questionnaires can be found in the supplementary material. To compensate participants, we offered a \$25 raffle reward for every 25 participants who completed the survey.
@@ IV-C Survey Results
Both stakeholders agreed that guidelines could be beneficial for creating better visualizations (Practitioners: 27%, Researchers: 28%), improving confidence in design decisions (P: 20%, R: 23%), and facilitating better communication with stakeholders (P: 20%, R: 22%). However, researchers weighed the usefulness of guidelines on increasing visualization design knowledge less (15%) than practitioners (24%). In the Other comment section, several practitioners mentioned additional benefits of guidelines: serving as a “standard of practice”; helping avoid “misleading, ambiguous”, and “deceitful” visualization designs; increasing “user understanding” and “trust”, producing “consistent user experience”; “facilitating design decision making”. Researchers’ comments were similar, while they additionally mentioned that guidelines provide a “scaffold” or “starting point” for learners.
@@ IV-E Interview Participants
TABLE I: Demographic characteristics of practitioner and researcher interviewees
@@ IV-F Interview Analysis
All interviews were conducted within a 30 to 45-minute timeframe, and each was transcribed using an automatic transcription service. The transcripts were then manually checked for errors. During the initial coding process, a research assistant pulled related quotes from the transcripts, excluding unrelated conversational text for efficient analysis. The assistant coded them in a spreadsheet using the interview questions as the initial set of codes. The lead researcher reviewed and refined the codes. They repeated this process for approximately one-third of the interviews until arriving at a consistent set of codes (i.e., inductive coding) . The research assistant then applied the final codes to the remaining interviews, which were comprehensively reviewed by the lead researcher (i.e., deductive coding). Overall, we categorized the quotes into eleven themes: five focused on experience and perception, challenges related to guidelines, two themes concerning perception and accessibility of empirical studies, four themes addressing the perception of the knowledge gap, strategies to mitigate this gap, and expectations within the visualization communities.
@@ IV-G2 Examples of Visualization Guidelines
Researchers echoed practitioners in some areas, reiterating guidelines on zero-baselines (R1, R10) and pie charts (R7). But they also shared more specific design guideline knowledge such as rainbow color maps (R2, R6, R9), length vs angle vs area (R1, R2, R10), comparison along a common-scale vs unaligned-scale (R1, R8), 3D charts (R4), and banking to 45 degrees (R9). Some of these guidelines were neither found on our guideline analysis nor discussed by practitioners. Researchers were often critical in their discussions, noting exceptions to general guidelines and pointing out vagueness in some common advice. For instance, R6 said, “there are certain cases when with the scientific data it makes sense to actually utilize [rainbow colormaps]”, while R7 said, “One vis design guideline that we hear a lot about is don’t use pie charts, which I would take to be a bit vague.”
@@ IV-I4 Ideas and Methods for Alleviating the Gap
V Takeaways
Overall, there was a fair degree of agreement between both stakeholders, with slight differences in emphasis. They agreed that guidelines can be valuable for providing knowledge and educational benefits; however, the current formats are not ideal. Specifically, guidelines are often presented in a prescriptive manner and fail to consider many experiential and contextual factors. Similarly, both groups acknowledged the usefulness of empirical research but expressed concerns about its lack of accessibility in practice. It was evident that research papers, in their current form, are not suitable for practitioners, indicating a need for translation efforts, such as summaries or example-based descriptions using both good and bad examples.
Despite these points of agreement, there were also differences between the two groups. Practitioners rarely refer to research papers, and the guidelines they mentioned, as well as how they accessed them, differed from those cited by researchers. This validates our decision to filter research papers in the guideline collection and corroborates the discrepancies identified in Section 3. Regarding knowledge gaps, researchers were more aware of these gaps, while practitioners appeared less conscious of them, likely due to being limited by the resources available to them. In addressing these gaps, more effective engagement, such as providing publicly accessible content and fostering collaborative efforts (e.g., workshops, courses), were suggested, though there was some skepticism about the feasibility due to potential challenges, such as a lack of incentives.
VI Limitations
This study has several limitations. Sampling bias may be present, as participants were recruited from specific communities, which might not fully capture the diversity of practitioners and researchers. The qualitative nature introduces subjectivity in interpretation, despite efforts to ensure consistent coding. Additionally, self-reporting bias could impact our findings, as participants may have presented themselves more favorably on topics such as knowledge gaps or their use of guidelines. Finally, the varied experiences of participants make it challenging to aggregate their opinions into a unified perspective, which may affect the consistency of the findings.
VII Discussion
@@
First of all, there appears to be a need to establish consistent vocabularies for describing data visualizations, including chart names and task types. This issue was observed during our examination of guidelines and was reiterated in interviews with practitioners; e.g., there is often a lack of clear distinction between task types, chart types, and data types in the categorization of visualizations [datatoviz, visvocab, datavizproject]. While researchers may require a more granular vocabulary, such as the task taxonomy proposed by Brehmer and Munzner [brehmer2013multi], it must be tailored to the level of practical needs and contexts. Previous research has similarly advocated for eliminating ambiguity in vocabulary related to chart junk [akbaba2021manifesto], while others have argued that ambiguity in both research and practice can have its own value [shukla2022negotiating].
A critical issue we observed is inconsistent design guidelines available in practice, as well as authoritative guidelines that do not acknowledge contextual factors. Interviewees often recognized this issue and suggested the need for improving current design guidelines with empirical evidence details. Empirical studies often contain mixed and conflicting findings, reflecting the diverse contexts in which these studies are conducted. However, the challenge lies in effectively communicating these complex nuances to practitioners.
@@
We also observed how practitioners appreciate intuition and experience in design practice. We need to consider how to incorporate such experiential knowledge. VisGuides [diehl2018visguides] is an example of a platform where practitioners can share their examples for discussion and feedback. We can also imagine connecting existing examples to guidelines to create a triadic model of guidelines, empirical evidence, and relevant examples. The examples can demonstrate how people applied or deviated from the guidelines and present their success and failure cases. Furthermore, these examples might inspire researchers to develop future practice-oriented studies, thereby building a feedback and feedforward loop.
@@
Participants also shared the difficulty of finding relevant guidelines. Therefore, annotating the guidelines with detailed metadata, such as tasks, chart types, or other contextual factors, would be necessary to facilitate efficient navigation of the landscape of visualization design guidelines. Current efforts by practitioners [datatoviz, datavisualizationcatalogue, datavizproject] provide hints on how this idea can be realized. For example, a decision tree can be used to find a chart based on data types or task goals. However, as mentioned earlier, practitioners’ taxonomy is currently not consistent and streamlined compared to the research side. To alleviate this issue, we could utilize templates to create more structured and easily consumable guidelines, as demonstrated in Figure 2. These templates can also be transformed into computational constraints to enhance a visualization recommendation system [moritz2018formalizing, zeng2023too].
@@
While building such a design guideline knowledge base is feasible given the relatively young field of visualization the substantial ongoing challenge lies in maintaining its relevance and functionality over time. We believe that it will require collective efforts from both practitioners and researchers. Even if there have been efforts for both stakeholders to meet and collaborate, we have not yet seen a cohesive community that brings both together. For example, IEEE VIS primarily caters to researchers, while the Data Visualization Society focuses on practitioners. Given the challenges of creating a new community site [kraut2012building], a viable alternative is to utilize existing communities to foster collaborations.
@@
Practitioners often write guidelines to promote their commercial tools (e.g., DataWrapper [datawrapper]) or attract clients (e.g., Data Viz Project [datavizproject]), similar to how researchers write blog articles for recognition. Yet, interviews highlighted a lack of incentives as a major issue. Addressing this, we should consider providing both intrinsic and extrinsic motivation to support the knowledge base’s maintenance and development. Ko’s market-based incentive model [reviewmodel], offering tokens for contributions, presents an intriguing solution. Additionally, harnessing recent advancements in large language models (LLM) could offer intelligent assistance in converting empirical results into practical guidelines [kim2023good]. For instance, LLMs can efficiently extract summary insights from research papers, translate these insights into practitioner-friendly language, and create relevant examples through code generation and contextual understanding.
@@
The field of data visualization is not alone in facing a knowledge gap between research and practice; this is a common challenge across many disciplines. In medicine, evidence-based practices integrate clinical expertise with the best available research evidence [sackett1996evidence]. In policy-making, there is an emphasis on translating research findings into actionable policies and practices [bogenschneider2011evidence]. In education, it is common to use educational research to improve teaching and learning practices [davies1999evidence]. Learning from these established fields can also inform how the visualization field should progress to promote evidence-based design practice, such as developing knowledge brokering roles [ward2009knowledge] and investing collaborative research-practice partnerships [coburn2016research]. The recent VisGuides workshop also provides several insights into bridging the knowledge gap, highlighting the risks of guidelines suppressing creative solutions [visguides2022] and the development of practice-oriented guidelines [parsons2022considering].
VIII Conclusion